Marshall Inc. recently hired your consulting firm to improve the company's performance. It has been highly profitable but has been experiencing cash shortages due to its high growth rate. As one part of your analysis, you want to determine the firm's cash conversion cycle. Using the following information and a 365 day year, what is the firm's present cash conversion cycle? Enter your answer rounded to two decimal places. For example, if your answer is 123.45% or 1.2345 then enter as 1.23 in the answer box. Average Inventory = $75,000.00

Annual Sales = $650,000.00

Annual Cost of Goods Sold = $375,000.00

Average Accounts Receivable = $170,000.00

Average Accounts Payable = $30,000.00
Answer #1
Inventory turnover = COGS/inventory
Inventory turnover = 375000/75000
Inventory turnover = 5
days of inventory on hand = number of days in a year/inventory turnover
days of inventory on hand = 365/5
days of inventory on hand = 73
Receivables turnover = Credit sales/receivables
Receivables turnover = 650000/170000
Receivables turnover = 3.82
days of sales outstanding = number of days in a year/receivables turnover
days of sales outstanding = 365/3.82
days of sales outstanding = 95.55
Accounts payables turnover = purchases/payables
Accounts payables turnover = 375000/30000
Accounts payables turnover = 12.5
days of payables outstanding = number of days in a year/accounts payable turnover
days of payables outstanding = 365/12.5
days of payables outstanding = 29.2
Operating cycle = days of sales outstanding + days of inventory on hand
Operating cycle = 95.55+73
Operating cycle = 168.55
Cash conversion cycle = Operating cycle - days of payables outstanding
Cash cycle = 168.55-29.2
Cash cycle = 139.35